find the measure of each angle in a triangle if the first angle measures 3 degrees more than the second angle which contains the same number of degrees as the third angle.
Answer by macston(5194)

.
x=second angle=third angle; x+3=first angle
Sum of angles in triangle = 180 degrees
first angle+second angle+third angle=180 degrees
(x+3)+x+x=180 degrees Subtract 3 from each side.
3x=177 degrees Divide each side by 3.
x=59 degrees
ANSWER 1: The second and third angles each measure 59 degrees.
x+3=59 degrees+3=62 degrees ANSWER 2: The first angle measures 62 degrees.